AIDA GELINA BRIKEN nToF CRIB ISOLDE CIRCE nTOFCapture DESPEC DTAS EDI_PSA 179Ta CARME StellarModelling DCF K40
  CARME  ELOG logo
Message ID: 301     Entry time: Sun Feb 27 10:00:07 2022
Author: RSS TD 
Subject: Current status 

Ring current ~9e7 particles/s

Merger, Tape server (WS3): 1414 kB/s writing ok

Leakage current (WS4): 3.43 uA

DAQ (WS2): temperature checked, clock status: ok, adc calibration: ok, white rabbit check: failed

Online monitoring (WS 5): done

Attachment 1: 11-xy-2022-02-27_11-01-22.png  271 kB  | Hide | Hide all
11-xy-2022-02-27_11-01-22.png
Attachment 2: 11-ey-rate2022-02-27_11-00-46.png  157 kB  | Hide | Hide all
11-ey-rate2022-02-27_11-00-46.png
Attachment 3: 11-ex-rate-2022-02-27_11-00-27.png  154 kB  | Hide | Hide all
11-ex-rate-2022-02-27_11-00-27.png
ELOG V3.1.4-unknown